GDL just paid a dividend of $4.06 per share. You expect dividends to grow 12% for the next 3 years, 10% the year after that, and then grow at 4% per year forever. If the required return is 14%, what is the price of the stock today? Round your answer to 2 decimal places, for example $10.12.
Answer #1

D1=(4.06*1.12)=4.5472

D2=(4.5472*1.12)=5.092864

D3=(5.092864*1.12)=5.70400768

D4=(5.70400768*1.1)=6.27440845

Value after year 4=(D4*Growth rate)/(Required return-Growth rate)

=(6.27440845*1.04)/(0.14-0.04)

=65.2538479

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.5472/1.14+5.092864/1.14^2+5.70400768/1.14^3+6.27440845/1.14^4+65.2538479/1.14^4

=$54.11(Approx)